The angles of a triangle form a ratio of 3:3:4. Find the measure of the largest angle.
Answer by Cromlix(4381)

First add all the ratios
= 10
Divide the total number of degrees
in a triangle = 180 degs.
by 10 = 18degs
Therefore your sizes are 3*18 = 54 degrees (twice)
and the largest angle is 4 * 18 = 72 degs.
Proof:
54 + 54 + 72 = 180 degrees.
Hope this helps
:-)